# Testing and Enhancing Multi-Agent Systems for Robust Code Generation

### v1

@@ Abstract.
Multi-agent systems (MASs) have emerged as a promising paradigm for automated code generation, demonstrating impressive performance on established benchmarks by decomposing complex coding tasks across specialized agents with different roles. Despite their prosperous development and adoption, their robustness remains pressingly under-explored, raising critical concerns for real-world deployment. This paper presents the first comprehensive study examining the robustness of MASs for code generation through a fuzzing-based testing approach. By designing a fuzzing pipeline incorporating semantic-preserving mutation operators and a novel fitness function, we assess mainstream MASs across multiple datasets and LLMs. Our findings reveal substantial robustness flaws of various popular MASs: they fail to solve 7.9%–83.3% of problems they initially resolved successfully after applying the semantic-preserving mutations. Through comprehensive failure analysis, we identify a common yet largely overlooked cause of the robustness issue: miscommunications between planning and coding agents, where plans lack sufficient detail and coding agents misinterpret intricate logic, aligning with the challenges inherent in a multi-stage information transformation process. Accordingly, we also propose a repairing method that encompasses multi-prompt generation and introduces a new monitor agent to address this issue. Evaluation shows that our repairing method effectively enhances the robustness of MASs by solving 40.0%–88.9% of identified failures. Re-execution of the fuzzing process on the repaired MASs shows that the number of found failures decreases up to 85.7%, demonstrating that repaired MASs exhibit superior robustness. Our work uncovers critical robustness flaws in MASs and provides effective mitigation strategies, contributing essential insights for developing more reliable MASs for code generation.
@@
conference: The ACM International Conference on the Foundations of Software Engineering; 5-9 July, 2026; Montreal, Canadaccs: Software and its engineering Software testing and repairingccs: Software and its engineering Robustness
@@ 1. Introduction
The advent of large language models (LLMs) has fundamentally transformed code generation methodologies, enabling the automated production of code snippets that fulfill user requirements expressed in natural language (Jiang et al., 2024; Chen et al., 2024a). Among these methodologies, multi-agent systems (MASs) have exhibited particular promise. Implemented through the same backend LLM, MASs decompose complex coding tasks into manageable sub-tasks and distribute them among specialized agents with distinct roles (Dong et al., 2024; Hong et al., 2023; Zhang et al., 2024a; Islam et al., 2024; M. A. Islam, M. E. Ali, and M. R. Parvez, 2025). For instance, planning agents are responsible for planning the whole coding logic, while coding agents focus on implementing the code based on the coding logic. Through the deployment of carefully crafted prompts and communication protocols, these specialized agents collaborate effectively, yielding impressive results across diverse coding challenges on the benchmark datasets (Chen et al., 2021b; Austin et al., 2021; Dong et al., 2025; Li et al., 2022; Yu et al., 2024b).
Despite the effectiveness of MASs for code generation across various benchmarks, their robustness remains pressingly under-explored. The lack of robustness in MASs can pose significant risks to real-world deployment. For instance, when deployed in production environments, non-robust MASs may generate inconsistent and incorrect code solutions, resulting in system failures and increased debugging costs (Mastropaolo et al., 2023). Moreover, the robustness issue in MASs can be further exacerbated by the inherent instability of LLMs and the variability of users’ natural language input (Wang et al., 2025).
@@
To understand the robustness characteristics of MASs, this paper conducts the first study on the robustness of MASs for code generation using a fuzzing-based approach. Inspired by existing efforts in the community on AI model robustness testing (Wang et al., 2023b; Mastropaolo et al., 2023; Chen et al., 2021a; Iakusheva et al., 2023; Xie et al., 2022; Rehman and Srinivasan, 2023), we design four semantic-preserving mutation operators over the input questions of MASs. The fuzzing is guided by a fitness function that evaluates the deviation of MAS-generated plans and codes to provide feedback for mutation. Our fuzzing results reveal serious robustness flaws: three mainstream MASs (Dong et al., 2024; Hong et al., 2023; Zhang et al., 2024a) fail to solve 7.9%–83.3% of questions that they initially resolved correctly.
Based on the fuzzing results, we summarize several key observations that depict the robustness flaws in contemporary MASs. In particular, we for the first time reveal that the communication gap between coding agents and planning agents (“planner-coder gap”) constitutes a major cause of MASs’ robustness flaws, accounting for an overwhelming 75.3% of all failures. Specifically, planning agents tend to generate coding plans that are logically correct but lack certain detailed explanations, and coding agents are prone to misinterpretations of specific expressions and complex logic. We uncover that this gap significantly compromises communication and understanding between agents. Through our analysis, we identify five error patterns that characterize this gap, revealing that the key to enhancing MAS robustness lies in bridging the planner-coder gap.
To address these issues, we propose a novel repairing method to enhance the robustness of MASs. As the first step, we present a novel formulation of the code generation paradigm of MASs and demonstrate that this multi-stage transformation process can introduce information loss and semantic drift across agent boundaries. To mitigate the information loss, our repairing method incorporates two primary components: multi-prompt generation and monitor agent insertion. Multi-prompt generation is designed to diversify the expressions in original input questions and reduce the chance that MASs misunderstand specific expressions. Specifically, we implement additional mutations on input questions to generate diverse prompts, providing the system with potentially better-formulated expressions that can reduce misinterpretation during agent communication. Moreover, we introduce a specialized monitor agent to bridge the planner-coder gap. This agent serves as an intermediary to mitigate the information loss between the planning agent and the coding agent, facilitating better coordination through two key processes: a plan interpretation process that enhances the coding agent’s comprehension by providing a detailed interpretation of received plans, and a code check process, where the monitor agent reviews and validates the alignment of the generated code to the interpreted plan.
@@
We conduct extensive experiments to evaluate our repairing method across three prominent MASs (Dong et al., 2024; Hong et al., 2023; Zhang et al., 2024a) with three backend LLMs (Achiam et al., 2023; Zhu et al., 2024; OpenAI, 2022) on four datasets (Dong et al., 2025; Li et al., 2022; Yu et al., 2024b). Results demonstrate that our approach can repair 40.0%–88.9% of failures identified during fuzzing. We also re-execute the fuzzing process on the repaired MASs and find that the number of failures decreases up to 85.7%, showing that repaired MASs exhibit superior robustness. Additionally, ablation studies demonstrate the necessity of both multi-prompt generation and monitor agent insertion to mitigate different robustness shortcomings of MASs. Overall, we summarize our contributions as follows:
@@
  • We advocate the first systematic study on the robustness of MASs for code generation, focusing on the impact of semantically equivalent requirements on MAS performance. Our fuzzing-based testing approach incorporates semantic-preserving mutations and a well-designed fitness function to effectively assess mainstream MASs. We summarize key findings on the ill-robustness of current MASs in this domain.
  • For the first time, we identify a major cause of robustness issues in MASs: the “planner-coder gap,” which compromises communication and understanding between agents. From our observation, we analyze the semantic drift across agent boundaries and formulate the code generation paradigm of MASs. Accordingly, we propose a novel repairing method, which encompasses multi-prompt generation and a new monitor agent, to mitigate the planner-coder gap.
  • We conduct comprehensive experiments validating our repairing method, demonstrating its effectiveness in augmenting the robustness of MASs to a large extent.
@@ 2. Background
Code generation has been a central focus of software engineering research, aiming to automate programming tasks and enhance developer productivity by generating desired code from natural language specifications (Chen et al., 2022; Shin and Nam, 2021; Dehaerne et al., 2022; Chen et al., 2025; Cao et al., 2024; Du et al., 2024; Mathews and Nagappan, 2024; Gao et al., 2024). MASs are artificial intelligence systems that leverage LLMs as their core computational engines to simulate domain experts with specialized capabilities (Qian et al., 2023; Park et al., 2023; Xu et al., 2023). With advances in prompt engineering (Liu et al., 2023), developing MASs to enhance code generation has emerged as a prominent research direction (He et al., 2024; Liu et al., 2024; Tran et al., 2025).
@@
Fig. 1 depicts a typical workflow of a code generation MAS. Upon receiving a natural language requirement, the system routes the input to the planning agent (planner) to generate a clear and correct coding plan. Note that the planning agent does not simply paraphrase the user input. Instead, it serves as an important coding step that decomposes the coding problem into a set of requirements and designs several key logical steps to solve the problem, which determines the algorithm and the logic for code implementation (Zhang et al., 2024a; Dong et al., 2024). The coding agent (coder) then implements the solution following the provided plan. Subsequently, the testing agent (tester) automatically generates test cases and evaluates the implementation, producing a test report. If the code fails, the report is returned to the coder for iterative refinement until the solution passes all tests or the computational budget is exhausted.
This specialized multi-agent workflow represents a fundamental paradigm shift from single-LLM code generation approaches. Unlike the end-to-end paradigm where a unified model directly transforms natural language into executable code (Chen et al., 2021b; Wang et al., 2021; Fried et al., 2023; Nijkamp et al., 2022), MASs decompose the complex task into specialized subtasks handled by domain-specific agents. This paradigm demonstrates remarkable advantages in handling complex coding tasks. Nevertheless, our implication in Sec. 5 reveals critical robustness issues concerning the interaction between planner and coder.
@@ 3. Motivation
MASs (Hong et al., 2023; Dong et al., 2024; Zhang et al., 2024a; Huang et al., 2023; Zhang et al., 2024b; Ishibashi and Nishimura, 2024) have emerged as a transformative technology with immense commercial potential. IT companies have put much effort into the development of MASs, with Google Cloud announcing comprehensive enhancements to Vertex AI that include the Agent Development Kit (ADK) and Agent Engine for building and deploying enterprise-grade MASs at scale (clo, 2025). OpenAI has also launched new APIs and tools specifically designed for agentic applications, including the Responses API and the Agents SDK for orchestrating multi-agent workflows, demonstrating growing commercial interest in MASs (ope, 2025).
However, assessing the reliability and robustness of these complex systems remains a blank. Unlike single agents and individual models, MASs involve intricate inter-agent communications, dynamic role assignments, and collaborative decision-making processes that make their behavior difficult to predict and evaluate systematically (He et al., 2024). This gap in assessment capabilities becomes particularly concerning given the high-stakes commercial applications these systems are designed to serve (Hong et al., 2023). Inspired by previous work that employs testing to assess the robustness of AI systems (Li et al., 2023; Mastropaolo et al., 2023; Xie et al., 2022), we propose a new way to gauge the robustness of MASs for code generation based on the semantically equivalent input mutations. In particular, we presume that a robust MAS should sustain the consistency when addressing semantically equivalent input and generating correct programs regardless of how developers articulate their requirements. This idea stems from the observation that developers’ requests typically vary in vocabulary and phrasing due to the variability of natural language.
To illustrate this idea, consider a simple example where a developer might request: (1) “return n-th Fibonacci number,” (2) “give back the n-th number in the Fibonacci sequence.” Although these requests use different phrasing, their semantic meaning is identical. A robust MAS should generate functionally correct solutions for both expressions. However, we find that even the state-of-the-art MASs suffer from inconsistencies in handling such semantically equivalent expressions. For example, PairCoder (Zhang et al., 2024a) generates correct code for the first request, while failing on the second due to incorrect edge case handling.
@@ 4. Fuzzing MAS with Semantically Equivalent Requirements
To fill the blank in understanding of MAS’s robustness in code generation as discussed in Sec. 3, in this work, we conduct the first fuzz testing with a framework specifically designed for MASs. Our approach systematically generates semantically equivalent variations of programming requirements, exhaustively testing MASs’ ability to handle diverse expressions of the same requirement (coding problem), with the goal of comprehending the robustness of MASs. In particular, this fuzzing framework is specifically designed to expose robustness flaws arising from the unique inter-agent communication issues in MASs using a dual-component fitness function, which, for the first time, assesses the impact of mutations on both the final code output and the intermediate planning artifacts. This section elaborates on the design of this fuzzing methodology, the setup of fuzzing experiments, and the fuzzing results.
@@ 4.1. Fuzzing Methodology
We first present four semantic-preserving mutation operators that generate alternative expressions of selected questions. Then, we introduce a fitness function that calculates rewards based on the differences in both output code and generated plans between the original and mutated questions, to provide comprehensive guidance for subsequent fuzzing iterations.
@@
Table 1. Semantic-preserving mutation operators and description.
@@
Table 1 presents these four mutation operators: Rephrase, Insert, Expand, and Condense. We implement these operators using GPT-4o (Achiam et al., 2023), a state-of-the-art LLM, through carefully designed prompts (available in our artifact [^1] https://github.com/anonymous-for-submit/MAS_testing_repairing/tree/main/More_results/Mutation_opertors). Our approach effectively addresses two key challenges from previous work (Wang et al., 2023b; Mastropaolo et al., 2023): (1) it avoids semantic deviations that may occur during mutation, and (2) it produces natural and fluent expressions in the mutated questions.
Following Yu et al. (Yu et al., 2024a), our operators cover prevalent semantic-preserving transformations, providing comprehensive coverage of semantic variations. Their preliminary work (Yu et al., 2024a) demonstrates that LLM-generated mutations effectively simulate diverse human expressions. Our manual verification confirms that the mutated questions produced by our operators maintain consistent human linguistic patterns while preserving the original semantics, making them suitable for testing the robustness of MASs.
@@
Code Reward. To mitigate the occasional failure due to randomness rather than underlying robustness issues, we execute MASs for $n$ times for each question. We collect the generated code and calculate the rate of completions passing all test cases provided in the dataset. Based on the insight that an increasing number of failure trials on the mutated questions indicate degraded MAS performance (Wang et al., 2023a; Chen et al., 2021b), we calculate the code reward by measuring the difference between the pass rates of code generated from original versus mutated questions:
@@
where $n$ represents the number of attempts, $c_{i}$ and $\hat{c}_{i}$ indicates whether the $i$-th generation passes the test (1 for pass, 0 for fail). Positive rewards indicate that the mutated question reduces MAS performance, making it valuable for identifying robustness issues in subsequent fuzzing processes.
Plan Reward. Apart from the generated code, we also compare the semantic of plans generated by the planning agent for the original and mutated questions. Significant changes in the generated plans often indicate potential robustness issues in the planning phase, which frequently make substantial influence over the coding agent’s ability and the final code quality. We apply Sentence-BERT (Reimers and Gurevych, 2019) to obtain the embedding and calculate semantic similarity between corresponding plans. The plan reward is computed by subtracting the normalized similarity from 1:
@@
The final fitness function is calculated as the sum of the code reward and the plan reward.
Input: Multi-Agent System $M$, Seed Pool $\mathcal{S}$, Mutation Operators $\mathcal{O}$, Fitness Function $\mathcal{F}$, Budget $\mathcal{B}$, Attempts $n$
@@
Output: Failure Questions Set $\mathcal{R}$
@@
4 $ConsumedQueries\leftarrow 0$;
5
6while $ConsumedQueries<\mathcal{B}$ and $\mathcal{S}\neq\emptyset$ do

### v2

# Understanding and Bridging the Planner-Coder Gap: A Systematic Study on the Robustness of Multi-Agent Systems for Code Generation

@@ Abstract.
Multi-agent systems (MASs) have emerged as a promising paradigm for automated code generation, demonstrating impressive performance on established benchmarks by decomposing complex coding tasks across specialized agents with different roles. Despite their prosperous development and adoption, the fundamental mechanisms underlying their robustness remain poorly understood, raising critical concerns for real-world deployment. This paper conducts a systematic empirical study to uncover the internal robustness flaws of MASs using an automated mutation-based methodology. By designing a testing pipeline incorporating semantic-preserving mutation operators and a novel fitness function, we assess mainstream MASs across multiple datasets and LLMs. Our findings reveal substantial robustness flaws: semantically equivalent inputs cause drastic performance drops, with MASs failing to solve 7.9%–83.3% of problems they initially resolved successfully after applying the semantic-preserving mutations.
Through comprehensive failure analysis, we discover a fundamental cause underlying these robustness issues: the planner-coder gap, which accounts for 75.3% of failures. This gap arises from information loss in the multi-stage transformation process where planning agents decompose requirements into underspecified plans, and coding agents subsequently misinterpret intricate logic during code generation. Based on this formulated information transformation process, we propose a repairing method that mitigates information loss through multi-prompt generation and introduces a monitor agent to bridge the planner-coder gap. Evaluation shows that our repairing method effectively enhances the robustness of MASs by solving 40.0%–88.9% of identified failures. Re-execution of the testing process on the repaired MASs shows that the number of detected failures decreases up to 85.7%, demonstrating that repaired MASs exhibit superior robustness. Our work uncovers critical robustness flaws in MASs and provides effective mitigation strategies, contributing essential insights for developing more reliable MASs for code generation.
@@
conference: The ACM International Conference on Software Testing and Analysis; 3-9 Oct, 2026; Oakland, California, United Statesccs: Software and its engineering Software testing and repairingccs: Software and its engineering Robustness
@@ 1. Introduction
The advent of large language models (LLMs) has fundamentally transformed code generation methodologies, enabling the automated production of code that fulfills user requirements expressed in natural language (Jiang et al., 2024a; Hou et al., 2024). Among these methodologies, multi-agent systems (MASs) have emerged as a particularly promising paradigm. By leveraging the same backend LLM, MASs decompose complex coding tasks into manageable sub-tasks and distribute them among specialized agents with distinct roles (Dong et al., 2024; Hong et al., 2024; Zhang et al., 2024a; Islam et al., 2024, 2025). For instance, planning agents design the overall coding logic and algorithmic structure, while coding agents implement the actual code based on these plans. Through carefully crafted prompts and communication protocols, these specialized agents collaborate effectively, achieving impressive performance across diverse coding challenges on established benchmark datasets (Chen et al., 2021a; Austin et al., 2021; Li et al., 2022; Yu et al., 2024a).
Despite their demonstrated effectiveness on benchmarks, a critical limitation exists in our understanding of MAS robustness for code generation. While MASs work well on standard evaluation based on test cases, the fundamental mechanisms underlying their failures remain poorly understood—particularly when faced with semantically equivalent variations of input requirements. This “black box” problem poses significant risks for real-world deployment: non-robust MASs may generate inconsistent or incorrect code solutions, resulting in system failures and increased debugging costs (Mastropaolo et al., 2023; Zhong et al., 2024). Without a systematic understanding of why and how MASs fail under semantic-preserving input variations, it remains difficult to develop principled solutions for enhancing their reliability.
@@
To address this knowledge gap, we conduct a systematic empirical study to uncover the internal robustness flaws of MASs using an automated mutation-based methodology. Inspired by existing efforts on AI model robustness testing (Wang et al., 2023a; Mastropaolo et al., 2023; Chen et al., 2021b; Iakusheva et al., 2023; Xie et al., 2022; Rehman and Srinivasan, 2023), we design four semantic-preserving mutation operators that systematically transform input questions while preserving their intended semantics. Our testing methodology is guided by a novel fitness function that evaluates both the quality of generated code and the semantic similarity of generated plans, providing fine-grained feedback for mutation selection. We assess three mainstream MASs (Dong et al., 2024; Hong et al., 2024; Zhang et al., 2024a) across multiple datasets and backend LLMs. Our empirical findings reveal substantial robustness flaws: semantically equivalent inputs cause drastic performance drops, with MASs failing to solve 7.9%–83.3% of problems they initially resolved successfully after applying semantic-preserving mutations.
Through comprehensive failure analysis of the testing results, we discover a fundamental cause underlying these robustness issues: the planner-coder gap, which accounts for 75.3% of all observed failures. This gap arises from information loss in the multi-stage transformation process inherent to MAS architectures. Specifically, planning agents decompose user requirements into high-level plans that are often logically sound but lack sufficient implementation details, while coding agents subsequently misinterpret intricate logic or ambiguous expressions during code generation. We systematically characterize this gap through five distinct error patterns (EPs): EP-1: Gap in Core Concepts, where coding agents misunderstand fundamental algorithmic concepts in plans; EP-2: Gap in Edge Cases, where boundary conditions specified in plans are overlooked during implementation; EP-3: Gap in Complex Logic, where multi-step reasoning or nested conditions are incorrectly translated; EP-4: Gap in Relational Phrases, where quantity or spatial relationships are misinterpreted; and EP-5: Gap in Condition Judgments, where conditional branches or guard clauses are improperly implemented. These patterns reveal that enhancing MAS robustness fundamentally requires bridging the planner-coder gap to mitigate information loss across agents.
Based on our formulated understanding of the information transformation process, we propose a repairing method to mitigate information loss and bridge the planner-coder gap. Our approach incorporates two synergistic components. First, multi-prompt generation diversifies input question expressions through semantic-preserving mutations, reducing the likelihood of misinterpretation. Second, we introduce a specialized monitor agent as an intermediary between planning and coding agents. This monitor agent operates through two key processes: (1) a plan interpretation process that provides detailed explanations to enhance the coding agent’s comprehension, and (2) a code validation process that verifies alignment between generated code and the interpreted plan. By explicitly addressing information loss at the critical planner-coder interface, our monitor agent effectively bridges the communication gap underlying most robustness failures.
@@
We conduct extensive experiments to evaluate our repairing method across three prominent MASs (Dong et al., 2024; Hong et al., 2024; Zhang et al., 2024a) with three backend LLMs (Achiam et al., 2023; Zhu et al., 2024; OpenAI, 2022) on four datasets (Chen et al., 2021a; Austin et al., 2021; Li et al., 2022; Yu et al., 2024a). Results demonstrate that our approach successfully repairs 40.0%–88.9% of identified failures. When we re-execute the testing process on repaired MASs, discovered failures decrease by up to 85.7%, providing empirical evidence of substantially improved robustness. Ablation studies confirm that both components are necessary, as they address complementary aspects of the robustness challenge. Overall, we summarize our contributions as follows:
@@
  • We conduct the first systematic empirical study on the robustness of MASs for code generation, focusing on their behavior under semantically equivalent input variations. Our automated mutation-based methodology incorporates semantic-preserving mutations and a novel fitness function to effectively assess mainstream MASs, revealing substantial robustness flaws where performance drops by 7.9%–83.3% under semantic-preserving transformations.
  • We discover and characterize the planner-coder gap as a fundamental cause of MAS robustness failures, accounting for 75.3% of observed issues. Through systematic analysis, we identify five distinct error patterns that manifest this gap and formulate the multi-stage information transformation process in MAS code generation, providing theoretical grounding for understanding information loss across agent boundaries.
  • Based on our empirical findings and theoretical formulation, we propose a repairing method that mitigates information loss through multi-prompt generation and monitor agent insertion. Comprehensive experiments demonstrate that our approach repairs 40.0%–88.9% of identified failures and reduces failure rates by up to 85.7% in subsequent testing, validating its effectiveness in enhancing MAS robustness.
@@ 2. Background
Code generation has been a central focus of software engineering research, aiming to automate programming tasks and enhance developer productivity by generating desired code from natural language specifications (Shin and Nam, 2021; Dehaerne et al., 2022; Cao et al., 2024; Du et al., 2024; Mathews and Nagappan, 2024; Gao et al., 2024). MASs are artificial intelligence systems that leverage LLMs as their core computational engines to simulate domain experts with specialized capabilities (Qian et al., 2024; Hong et al., 2024; Zhang et al., 2024a). With advances in prompt engineering (Liu et al., 2023), developing MASs to enhance code generation has emerged as a prominent research direction (He et al., 2024; Tran et al., 2025).
@@
Fig. 1 depicts a typical workflow of a code generation MAS. Upon receiving a natural language requirement, the system routes the input to the planning agent (planner) to generate a clear and structured coding plan. Note that the planning agent does not simply paraphrase the user input. Instead, it serves as an important coding step that decomposes the coding problem into a set of requirements and designs several key logical steps to solve the problem, which determines the algorithm and the logic for code implementation (Zhang et al., 2024a; Dong et al., 2024). The coding agent (coder) then implements the solution following the provided plan. Subsequently, the testing agent (tester) automatically generates test cases and evaluates the implementation, producing a test report. If the code fails, the report is returned to the coder for iterative refinement until the solution passes all tests or the computational budget is exhausted.
This specialized multi-agent workflow represents a fundamental paradigm shift from single-LLM code generation approaches. Unlike the end-to-end paradigm where a unified model directly transforms natural language into executable code (Wang et al., 2021; Fried et al., 2023), MASs decompose the complex task into specialized subtasks handled by domain-specific agents. This paradigm demonstrates remarkable advantages in handling complex coding tasks. Nevertheless, our analysis in Sec. 5 reveals critical robustness issues concerning the interaction between planner and coder.
@@ 3. Motivation
MASs (Hong et al., 2024; Dong et al., 2024; Zhang et al., 2024a) have emerged as a transformative technology with immense commercial potential. IT companies have invested significant effort in the development of MASs, with Google Cloud announcing comprehensive enhancements to Vertex AI that include the Agent Development Kit (ADK) and Agent Engine for building and deploying enterprise-grade MASs at scale (50). OpenAI has also launched new APIs and tools specifically designed for agentic applications, including the Responses API and the Agents SDK for orchestrating multi-agent workflows, demonstrating growing commercial interest in MASs (38).
However, assessing the reliability and robustness of these complex systems remains a significant research gap. Unlike single agents and individual models, MASs involve intricate inter-agent communications, dynamic role assignments, and collaborative decision-making processes that make their behavior difficult to predict and evaluate systematically (He et al., 2024). This gap in assessment capabilities becomes particularly concerning given the high-stakes commercial applications these systems are designed to serve (He et al., 2024). Inspired by previous work that employs testing to assess the robustness of AI systems (Li et al., 2023; Mastropaolo et al., 2023), we assess the robustness of MASs for code generation based on semantically equivalent input mutations. In particular, we presume that a robust MAS should maintain consistency when addressing semantically equivalent input and generating correct programs regardless of how developers articulate their requirements. This idea stems from the observation that developers’ requests typically vary in vocabulary and phrasing due to the variability of natural language.
To illustrate this idea, consider a simple example where a developer might request: (1) “return the n-th Fibonacci number”, (2) “give back the n-th number in the Fibonacci sequence.” Although these requests use different phrasing, their semantic meaning is identical. A robust MAS should generate functionally correct solutions for both expressions. However, we find that even the state-of-the-art MASs suffer from inconsistencies in handling such semantically equivalent expressions. For example, PairCoder (Zhang et al., 2024a) generates correct code for the first request, while failing on the second due to incorrect edge case handling.
@@ 4. Fuzzing MAS with Semantically Equivalent Requirements
To fill the gap in understanding of MASs’ robustness in code generation as discussed in Sec. 3, we design a fuzz testing framework specifically adapted to MAS architectures. Our approach systematically generates semantically equivalent variations of programming requirements to expose robustness flaws arising from inter-agent communication—a vulnerability dimension not addressed by traditional testing approaches. A key innovation is our dual-component fitness function, which evaluates mutations along two dimensions: (1) final code quality degradation and (2) semantic drift in intermediate plans. This MAS-specific design enables us to systematically uncover the planner-coder gap, a fundamental structural flaw that represents the dominant failure mode in contemporary MASs (see Sec. 5).
@@ 4.1. Fuzzing Methodology
We first present four semantic-preserving mutation operators that generate alternative expressions of selected questions. Then, we introduce a novel fitness function that calculates rewards based on the differences in both output code and generated plans between the original and mutated questions, to provide comprehensive guidance for subsequent fuzzing iterations.
@@
Table 1. Semantic-preserving mutation operators and descriptions.
@@
Table 1 presents these four mutation operators: Rephrase, Insert, Expand, and Condense. We implement these operators using GPT-4o (Achiam et al., 2023), a state-of-the-art LLM, through carefully designed prompts. This design effectively addresses two key challenges from previous work (Wang et al., 2023a; Mastropaolo et al., 2023): (1) it minimizes semantic deviations that may occur during rule-based mutation, and (2) it produces natural and fluent expressions in the mutated questions. Moreover, as shown in prior work (Yu et al., 2024b), LLM-generated mutations effectively simulate diverse human expressions. Our operators cover prevalent semantic-preserving transformations (Yu et al., 2024b), providing comprehensive coverage of semantic variations.
To confirm the reliability of our mutation operators in terms of semantic preservation, we conduct manual verification on 500 randomly sampled mutants, confirming that 99.2% of the mutants preserve semantics. Additionally, we measure the semantic similarity of all mutants to the original questions using Sentence-Transformers (Reimers and Gurevych, 2019), whose results show 98.4% average similarity.
@@
Code Reward. To mitigate the occasional failure due to randomness rather than underlying robustness issues, we execute MASs for $n$ times for each question. We collect the generated code and calculate the rate of completions passing all test cases provided in the dataset. Based on the insight that an increasing number of failure trials on the mutated questions indicate degraded MAS performance (Wang et al., 2023b; Chen et al., 2021a), we calculate the code reward by measuring the difference between the pass rates of code generated from original versus mutated questions:
@@
where $n$ represents the number of attempts, $c_{i}$ and $\hat{c}_{i}$ indicate whether the $i$-th generation passes the test (1 for pass, 0 for fail). Positive rewards indicate that the mutated question reduces MAS performance, making it valuable for identifying robustness issues in subsequent fuzzing processes.
Plan Reward. Apart from the generated code, we also compare the semantic of plans generated by the planning agent for the original and mutated questions. Significant changes in the generated plans often indicate potential robustness issues in the planning phase, which substantially influence the coding agent’s ability and the final code quality. We apply Sentence-Transformers (Reimers and Gurevych, 2019) to obtain the embedding and calculate semantic similarity between corresponding plans. The plan reward is computed by subtracting the normalized similarity from 1:
@@
The final fitness function is calculated as the sum of the code reward and the plan reward. This combination enables measuring robustness issues in both the code quality degradation and semantic drift in the planning stage, which is the key to comprehensively understanding MAS robustness.
Input: Multi-Agent System $M$, Seed Pool $\mathcal{S}$, Fitness Function $\mathcal{F}$, Budget $\mathcal{B}$, Attempts $n$
@@
Output: Failure Question Set $\mathcal{R}$
@@
4
5while $ConsumedQueries<\mathcal{B}$ and $\mathcal{S}\neq\emptyset$ do